The angle bisector of one angle of a triangle forms two angles that measure 27 degrees. The angle bisector of another angle forms two angles that measure 34 degrees. What is the measure of the third angle of the triangle?
Answer by dabanfield(803)

The angle bisector of one angle of a triangle forms two angles that measure 27 degrees. The angle bisector of another angle forms two angles that measure 34 degrees. What is the measure of the third angle of the triangle?
Since the bisectors divide the angles into two equal parts we know that two of the angles are 2*27 = 54 and 2*34= 68 degrees.
Since the angles of a triangle sum up to 180 we know then that for the third angle, x, we have:
x + 54 + 68 = 180
Solve the above for x.